If sin a + sin b = 2 , then find the value of cos a + cos b

Grade:10

1 Answers

Shaswata Biswas
7 years ago
We know -1\leqslant \sin a\leqslant 1, and -1\leqslant \sin b\leqslant 1
Here, sina + sinb = 2
It is only possible when sina = sinb = 1.
=> a = b = π/2
So the value of cosa + cosb = cos(π/2) + cos (π/2) = 0.
THANKS